Find the critical value (or values) for the t test for the given case n=12, a=0.01, left tailed

Find the critical value (or values) for the t test for the given case n=17, a=0.02, two tailed
Answer #1

1)

df = 12 - 1 = 11

\alpha = 0.01

t 0.01,11 = -2.718

Critical value = -2.718

2)

df = 17 - 1 = 16

\alpha / 2 = 0.02

t 0.01,16 = -2.718

Critical value = 2.583